### v2

# ,, thanks: These two authors contributed equally to this work thanks: These two authors contributed equally to this work

Abstract
The quantum Hall effect (QHE) is a cornerstone of topological physics, inspiring extensive explorations of its high-dimensional generalizations such as experimental realizations in synthetic systems including cold atoms, photonic lattices and metamaterials. However, the many-body effect in the higher dimensional QHE system remains poorly understood. In this study, we explore this problem by formulating the microscopic wavefunctions on a four dimensional sphere inspired by Laughlin’s seminal work. Employing a generalized pseudo-potential framework, we derive an exact microscopic Hamiltonian consisting of two-body projectors that annihilate the microscopic wavefunctions. Diagonalizations on finite system sizes show that the quasi-hole states remain zero energy while the quasi-particle states exhibit a finite gap, in consistency with an incompressible state. Furthermore, the pair distribution is calculated to substantiate the liquid-like nature of the wavefunction. Our work provides a preliminary understanding to the fractional topological states in high dimension.
The discovery of the integer quantum Hall effect (IQHE) Klitzing et al. (1980) marked a pivotal milestone in condensed matter physics research, opening up an era for studying topological phenomena. In IQHE systems, two-dimensional (2D) electrons subjected to a magnetic field form Landau levels, and the quantization of Hall conductance is intrinsically tied to Landau level wavefunction topologies Thouless et al. (1982). Moreover, many-body properties become sharp research focus following the discovery of the fractional quantum Hall effect (FQHE) Tsui et al. (1982). The groundbreaking insight comes from the celebrated Laughlin wavefunction Laughlin (1983), which captures the essence of the FQHE, including fractional Hall conductance and fractionally charged excitations. Various generalizations of the Laughlin wavefunction have been proposed, including the Read-Moore state Moore and Read (1991) and the Haldane-Rezayi state Haldane and Rezayi (1988), offering new perspectives on the FQHE.
Various investigations have explored the extension of Landau levels to higher-dimensional systems Zhang (2001); Karabali and Nair (2002); Bernevig et al. (2003); Jellal (2005); Li et al. (2012); Li and Wu (2013); Li et al. (2013); Peng et al. (2025).Most notably, Landau levels on the four-dimensional (4D) sphere ($S^{4}$) have been constructed Zhang (2001) by introducing an SU(2) Yang monopole located at the origin Yang (1978a), analogous to Haldane’s construction of Landau levels on $S^{2}$ with Dirac monopole Haldane (1983). The 4D quantum Hall effect (QHE) has been applied to the study of noncommutative geometry Chen et al. (2002); Susskind (2001); Hou and Peng (2002); Zhang (2003); Hasebe (2010, 2018, 2020, 2022, 2023).Experimental advances have also demonstrated the realization of the 4D QHE in cold atom systems Bouhiron et al. (2024); Price et al. (2015), optical lattices Lohse et al. (2018); Zilberberg et al. (2018), acoustic lattices Chen et al. (2021a, b); Xu et al. (2022); Chen et al. (2023a); Lai et al. (2024), and electric circuitsPrice (2020); Wang et al. (2020); Yu et al. (2020); Zhang et al. (2020). Furthermore, the Yang monopole has been observed in cold atom simulations and metamaterials Sugawa et al. (2018); Ma et al. (2021, 2023).Nevertheless, the many-body physics of the 4D FQHE remains poorly understood, constituting an exceptionally difficult and largely unexplored frontier that may be accessible through synthetic dimensions in future experiments.Yang et al. (2024); Ozawa and Price (2019); Wu et al. (2019); Guo et al. (2021); Jiang et al. (2021); Ezawa (2021); Lustig and Segev (2021); Kanungo et al. (2022); Trautmann et al. (2024).
In this article, we explore the many-body properties based on $S^{4}$ under the SU(2) monopole background. Both the Laughlin-type many-body wavefunctions as a power of a determinant type and the Jastrow type are constructed, which are ground states of Hamiltonians with suitably designed Haldane pseudo-potentials. Exact diagonalizations are performed on the corresponding Hamiltonians, revealing the signature of an incompressible quantum liquid state. These states provide further insights for studying high dimensional topological states.
Landau levels on an $S^{4}$ sphere arises from the quantization of a particle motion in the presence of the $SU(2)$ Yang monopole Yang (1978a). The single-particle Hamiltonian reads as Zhang (2001),
$H=\frac{(P_{\mu}+A_{\mu})^{2}}{2M},$ (1)
where the particle is confined on an $S^{4}$ sphere with radius $R$; $A_{\mu}=-\frac{1}{2R(R+x_{5})}\eta_{\mu\nu}^{a}I_{a}x_{\nu}$ is the $SU(2)$ gauge field of a Yang monopole, with the t’Hooft symbol $\eta_{\mu\nu}^{a}(\mu,\nu=1,2,3,4;a=1,2,3)$ ’t Hooft (1976); $I_{a}$ is the SU(2) generator with the Casimir $I(I+1)$ and $I$ the SU(2) quantum number. The single-particle wavefunctions to Eq.(1) are solved by the SO(5){SO(5)} irreducible representations (IRREP) $(N+2I,N)_{so5}$Yang (1978b, a); Zhang (2001), where $N\geq 0$ is the Landau level index. The eigen-energy is given $SO(5)$ Casimir $\mathcal{C}(p,q)$,
$E_{N}=\frac{\hbar^{2}}{2MR^{2}}(\mathcal{C}(N+2I,N)-2I(I+1))$ (2)
The lowest Landau level (LLL) states correspond to the IRREPs of $(2I,0)_{so5}$. Their degeneracy is
$d(2I,0)=\frac{1}{3!}(2I+1)(2I+2)(2I+3).$ (3)
We decompose the LLL according to the IRREPS of the subgroup $SO(4)$, with 6 generators can be constructed with t’Hooft symbol 111 The $\hat{J}_{a},\hat{K}_{a}$ satisfy commutation that $\hat{J}_{a},\hat{J}_{b}]=\text{i}\varepsilon_{abc}\hat{J}_{c},[\hat{K}_{a},\hat{K}_{b}]=\text{i}\varepsilon_{abc}\hat{K}_{c},[\hat{J}_{a},\hat{K}_{b}]=0$, see more detail in S.M.sup (2020)
$\hat{J}_{a}=\frac{1}{4}\eta^{a}_{\mu\nu}L_{\mu\nu},\ \ \,\hat{K}_{a}=\frac{1}{4}\bar{\eta}_{\mu\nu}^{a}L_{\mu\nu}.$ (4)
Applying the branching rule to the LLL states, we obtain $(2I,0)_{so5}=\bigoplus\limits_{j+k=I}(j,k)_{so4}$. Since $\hat{J}_{a}$ and $\hat{K}_{a}$ form two mutually commuting $SU(2)$ algebras, $(j,k)_{so(4)}$ denotes the direct product of two $SU(2)$ representations characterized by quantum numbers $j$ and $k$, respectively.
Next we explain how to express the LLL wavefunctions by the homogeneous polynomials of the 4-component fundamental $SO(5)$ spinors $\psi_{\alpha}$ with $\alpha=1\sim 4$. Through the second Hopf map, one can define a unit vector $x_{a}=\psi^{\dagger}_{\alpha}\Gamma^{a}_{\alpha\beta}\psi_{\beta},(a=1,\dots,5)$ on $S^{4}$. The $4\times 4$ Clifford algebra matrices $\Gamma^{a}$ mutually anticommute and their explicit representation is presented in S.M.Sec.II. Conversely, $\psi_{\alpha}$’s can be expressed through $x_{a}$ and a two-component spinor $u=(u_{1},u_{2})^{T}$ ,defining the isospin coordinates on $S^{2}$ via the 1st Hopf map ${n}_{i}=u^{\dagger}{\sigma}_{i}u$. (see S.M.Sec. II). The LLL wavefunctions can be organized into a $(2I+1)$-component spinor, with each component being an eigenstate of $I_{3}$, denoted by $f_{I_{3}}(x_{a})|II_{3}\rangle$, where $I\geq I_{3}\geq-I$. For compactness, $|II_{3}\rangle$ is represented on $S^{2}$ by $u_{1}^{I+I_{3}}u_{2}^{I-I_{3}}/\sqrt{(I+I_{3})!(I-I_{3})!}$. Then the LLL wavefunctions are expressed as
$f_{j,m_{1};k,m_{2}}^{(2I,0)}(x_{a},n_{i})={\cal N}\psi_{1}^{j+m_{1}}\psi_{2}^{j-m_{1}}\psi_{3}^{k+m_{2}}\psi_{4}^{k-m_{2}},$ (5)
where ${\cal N}=1/\sqrt{{(j+m_{1})!(j-m_{1})!(k+m_{2})!(k-m_{2})!}}$ is the normalization factor .
For later convenience in exploring many-body wavefunctions, we first construct the two-body states under the $SO(5)$ rotational symmetry on $S^{4}$, which is significantly more complex than the $S^{2}$ case with $SO(3)$ symmetry. The product of two single-particle LLL states is then decomposed into different $SO(5)$ channels. According to the branching rule ,
$(2I,0)_{so5}\otimes(2I,0)_{so5}=\bigoplus\limits_{n_{1},n_{2}}(n_{1}+n_{2},n_{1}-n_{2})_{so5},$ (6)
where $0\leq n_{1}\leq 2I$, and $0\leq n_{2}\leq n_{1}$. The simplest set of the two-body states lie in the channel of $(4I-2n,0)_{so5}$ with $0\leq n\leq 2I$. An SO(5) invariant is constructed as $S(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})\mathcal{R}_{\alpha\beta}\psi_{\beta}(x^{\prime}_{a},n_{i}^{\prime})$ with $\mathcal{R}$ being the charge conjugation matrix for the $SO(5)$ group , satisfying $\mathcal{R}^{2}=-1$, $\mathcal{R}^{\mathbf{T}}=\mathcal{R}^{-1}=\mathcal{R}^{\dagger}=$-$\mathcal{R}$  [^2] The explicit definition of $\mathcal{R}$ is shown in S.M. Sec. III. The $\mathcal{R}$ is defined as $\mathcal{R}=-\text{i}\Gamma_{3}\Gamma_{1}$. With the help of $\mathcal{R}$, we obtain the two-body states in the form
$\Phi_{J,M_{1};K,M_{2}}^{(4I-2n,0)}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\mathcal{N}_{J,M_{1};K,M_{2}}^{(4I-2n,0)}S^{n}$
$\times f_{J_{1},M_{1};J_{2},M_{2}}^{4I-2n}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime}),$ (7)
where $\mathcal{N}_{J,M_{1};K,M_{2}}^{(4I-2n,0)}$ is the normalization coefficient that is given by $SO(5)$ isoscalar (see S.M.Sec.VI), and $f_{J_{1},M_{1};J_{2},M_{2}}^{4I-2n}$ is a symmetric homogeneous polynomial with the degree of 4I2n4I-2n defined as
$f_{J_{1},M_{1};J_{2},M_{2}}^{4I-2n}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=C_{j_{1},m_{1};j_{1}^{\prime},m_{1}^{\prime}}^{J_{1},M_{1}}C_{j_{2},m_{2};j_{2}^{\prime},m_{2}^{\prime}}^{K,M_{2}}$
$\times f^{(2I-n,0)}_{j_{1},m_{1};k_{1},m_{2}}(x_{a},n_{i})f^{(2I-n,0)}_{j_{1}^{\prime},m^{\prime}_{1};k_{1}^{\prime},m_{2}^{\prime}}(x^{\prime}_{a},n_{i}^{\prime}),$ (8)
where $C_{j_{1},m_{1};j_{1}^{\prime},m_{1}^{\prime}}^{J_{1},M_{1}}$ is the SU(2) group Celebsh-Gordon coefficient . The statistics of the two-body wavefunciton is captured by the power of $S^{n}$: Odd and even values of $n$ correspond to the fermionic and bosonic statistics, respectively.
We move forward to the two-body states in more complex channels with $n_{1}>n_{2}$. Let us take the case of $n_{1}-n_{2}=1$ as an example which involves the regular $SO(5)$ harmonics in the absence of Yang monopole. In analogy to the SO(5) invariant $S$, an $SO(5)$ vector, defined as $X_{a}(x_{a},n_{i};x_{a}^{\prime},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})(\mathcal{R}\Gamma_{a})_{\alpha\beta}\psi_{\beta}(x_{a}^{\prime},n_{i}^{\prime})$, transforms under the IRREP $(1,1)_{so5}$ . $X_{a}$ can be decomposed into an SO(4) scalar $X_{5}$ and a 4-vector $X_{1\sim 4}$ transforming as $(\frac{1}{2};\frac{1}{2})_{so4}$(see S.M.sup (2020).Table(S3)). $S$ and $X_{5}$ are orthogonal to each other, since the former is SO(5) invariant and the latter is a component of five-vector. According to group theoretical analysis, the SO(5) of IRREP $(4I-2n-1,1)_{so5}$ with $0\leq n\leq 2I-1$ can be decomposed into two branches of SO(4) IRREPs as follows,
$(4I-2n-1,1)_{so5}$ (9)
$=$ $\bigoplus\limits_{j_{1}+j_{2}=2I-n-1}[(j_{1},j_{2})_{so4}\oplus(j_{1}+\frac{1}{2},j_{2}+\frac{1}{2})_{so4}].$
Hence, we have two sets of the two-body wavefunctions as
$\Phi_{J,M_{1};K,M_{2}}^{(4I-2n-1,1)}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\mathcal{N}_{J,M_{1};K,M_{2}}^{(4I-2n-1,1)}X_{5}S^{n}f_{J,M_{1},K,M_{2}}^{4I-2n-2}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime}),$
$\Phi_{J+\frac{1}{2},M_{1};K+\frac{1}{2},M_{2}}^{(4I-2n-1,1)}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\mathcal{N}_{J+\frac{1}{2},M_{1};K+\frac{1}{2},M_{2}}^{(4I-2n-1,1)}C_{\frac{1}{2},\alpha;J,m_{1}}^{J+\frac{1}{2},M_{1}}C_{\frac{1}{2},\beta;K,m_{2}}^{K+\frac{1}{2},M_{2}}\Psi_{\frac{1}{2},\alpha;\frac{1}{2},\beta}^{(1,1)}S^{n}f_{J,m_{1},K,m_{2}}^{4I-2n-2}(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime}),$ (10)
in which $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\pm\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{1}\pm iX_{2})$, and $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\mp\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{3}\pm iX_{4}).$ More generally, for $(n_{1}+n_{2},n_{1}-n_{2})_{so(5)}$, the two-body wavefunctions in these irreps are obtained using higher-rank $SO(5)$ harmonics, as detailed in S. M. Sec(IV).
To implement numerical calculations, we generalize the pseudo-potential formalism Haldane (1983); Jain (2007) from $S^{2}$ to $S^{4}$. The pseudo-potential, based on the two-body states constructed above, is represented by a projection operator as follows
$H=\sum\limits_{n_{1},n_{2}}V_{n_{1},n_{2}}P_{n_{1},n_{2}}$ (11)
[…]
We delve into the many-body microscopic wavefunctions. There are two different approaches to construct Laughlin wavefunction, namely the determinant-type and Jastrow-type Laughlin wavefunctions, respectively Laughlin (1983); Haldane (1983); Zhang (2001). It is noteworthy that the determinant-type one (Eq.(12) below) is not identical to Jastrow-type one (Eq.(17) below). The difference has been highlighted in the study of the FQHE states on the $\mathbb{CP}^{2}$ manifold Wang et al. (2023); Chern and Lee (2007). Nevertheless, both types of wavefunctions coincide on $S^{2}$.
The determinant-type Laughlin wavefunction takes the form of the Slater determinant raised to the $m$-th power.
[…]
where $N$ is the particle number, $A_{h}$ is the abbreviation of $SO(4)$ IRREP $(j_{h},m_{h,1};k_{h},m_{h,2})$ for the $h$-th particle,and $P$ is permutation from 1 to $N$. The filling of Eq.(12) is $\nu=\frac{d(2I,0)}{d(2mI,0)}$ , which approaches $1/m^{3}$ in the large-$m$ limit.
We focus on the determinant Laughlin wavefunction with $m=3$. For convenience, consider the case with the smallest Hilbert space. The determinant wavefunction of a full-shell with $I=\frac{1}{2}$, i.e., $N=4$, is $\Psi_{\mbox{det}}=\varepsilon_{P_{1}P_{2}P_{3}P_{4}}\Pi_{k=1}^{4}\psi_{k}(x_{P_{k},a},n_{P_{k},i})$. The determinant Laughlin state is defined as
$\Psi^{(m=3)}_{\mbox{det},N=4}=\left(\varepsilon_{P_{1}P_{2}P_{3}P_{4}}\prod\limits_{j=1}^{4}\psi_{j}(x_{P_{j},a},n_{P_{j},i})\right)^{3}$ (13)
By counting the power of SO(5) spinors, the single-particle LLL orbitals involved in Eq.(13) correspond to the case of $mI=\frac{3}{2}$ with the degeneracy $d(3,0)=20$. The full-shell determinant state is $SO(5)$ invariant; its power, the Laughlin determinant state, is likewise an $SO(5)$ singlet. Upon exchanging two particles, the wavefunction acquires a phase factor $(-1)^{3}=-1$, which ensures Fermi statistics. Hence, when expanding Eq.(13), the relative wavefunction between a pair of particles $k$ and $l$ only contains
$\left(\mathbf{\Psi}(x_{k,a},n_{k,i})\mathcal{M}\mathbf{\Psi}(x_{l,a},n_{l,i})\right)^{3},$ (14)
where the anti-symmetry matrix kernel $\mathcal{M}$ takes either $\mathcal{R}$ or $\mathcal{R}\Gamma^{a}$. Each anti-symmetric combination between two particles is denoted as one contraction. For example, the structure of Eq.(14) exhibits 3 contractions.
Figure 1: The energy spectra in the blocks marked by quantum numbers of $(M_{J},M_{K})$ for the Hamiltonian Eq.(15) with $N=4$ particles filling in the LLL orbital with the SO(5) IRREP of $(\frac{3}{2},0)_{so5}$. The two datasets (a) and (b) correspond to the sector $(0,0)$ and $(\frac{1}{2},\frac{1}{2})$,respectively. The orbital number $N_{o}=20$. The energy unit is taken as $\hbar^{2}/(2MR^{2})$, and the horizontal axis shows the index of states. The zero energy ground state is an $SO(5)$ singlet and unique. The lowest excitation is another $SO(5)$ single state with a gap $\Delta=4.8$.
Figure 2: The energy spectra for the same system as in Fig.(1): ($a$) quasi-hole states with $N=3$, ($b$) quasi-particle states with $N=5$. Only the block of $(\frac{1}{2},0)$ is represented. The ground states of the quasi-hole case remain at zero energy, while those of the quasi-particle case exhibit a finite gap. The ground states of both quasi-hole and quasi-particle cases are of the 20-dimensional SO(5) IRREP of $(3,0)_{so5}$.
Now we construct the pseudo-potential Hamiltonian for the many-body wavefunctions Eq.(12). The pseudo-potential Hamiltonian, constructed from two-body projection operators, is positive-definite and annihilates the wavefunction (12). Consequently, the annihilated wavefunction is a zero-energy ground state. For concreteness, the special case of Eq.(13) with $m=3,N=4$ is employed as an example, i.e.. Consider two fermions in the LLL orbits with $I^{\prime}=mI=\frac{3}{2}$. The two-body states satisfying the fermionic statistics are decomposed into SO(5) IRREPs of $(4I^{\prime}-2r-1-s,2r+1-s)_{so5}$ with $0\leq r<I^{\prime}$ and $0\leq s\leq 2r+1$. According to the two-body wavefunctions given in Eq.(VI.8) in S. M. Sec. (VI), the contraction number equals $2r+1$. Hence, the $r=0$ irreps, (4,0)SO(5)(4,0){SO(5)} and (5,1)SO(5)(5,1){SO(5)}, are orthogonal to the two-body sectors of the wavefunction in Eq. (13), as revealed by the contraction numbers. The above analysis can be easily generalize to an arbitary value of $N$ at $m=3$. Then the pseudo-potential Hamiltonian is constructed as
$H$ $=\sum\limits_{1\leq k<l\leq N}\Big{\{}V_{(4mI-1,1)}P_{4mI-1,1}(k,l)$
$+V_{(4mI-2,0)}P_{(4mI-2,0)}(k,l)\Big{\}},$ (15)
where $P_{m,n}(k,l)$ is the projection operator of two-particle states of the $k$-th and $l$-th ones.
Next we present the exact digonalization results for the case of $m=3$ and $N=4$. The $SO(5)$ is a rank-2 Lie group with two commutable generators, which can be taken as the sum of the diagonal operators of two inter-commutable $SU(2)$ subgroups defined in Eq.(4) of every particle: $J_{3}=\sum_{l=1}^{N}J_{3}(l)$ and $K_{3}=\sum_{l=1}^{N}K_{3}(l)$. The pseudo-potential Hamiltonian is diagonalized in blocks labeled by different values of $(M_{J},M_{K})$, which are $J_{3}$ and $K_{3}$ eigenvalues, respectively. Previously analysis shows that the wavefunction of Eq.(13) is an SO(5) singlet and a ground state of the Hamiltonian Eq. (15). Diagonalizations are performed and the ground state is found non-degenerate, hence, the wavefunction of Eq.(13) is the unique ground state. Energy spectra in the blocks of $(0,0)$ and $(\frac{1}{2},\frac{1}{2})$ with spectra shown in Fig.(1). The energies of the lowest and next lowest excited states are very close at $4.80$ and $4.93$ under the unit of $\frac{\hbar^{2}}{2MR^{2}}$. The lowest excited state is also an $SO(5)$ scalar as shown in the sector of $(0,0)$, and the next lowest excited states are 35-fold degenerate forming the SO(5) IRREP of $(4,0)_{{so5}}$. Such an $SO(5)$ IRREP is decomposed into different $SO(4)$ IRREPs carrying integer quantum numbers of $(2,0)_{so4}\oplus(1,1)_{so4}\oplus(0,2)$ with dimensions of $5$, 99, and $5$, respectively, and those carrying half-integer quantum numbers of $(\frac{3}{2},\frac{1}{2})_{so4}\oplus(\frac{1}{2},\frac{3}{2})_{so4}$ with dimensions 8 and 8, respectively, as shown in Fig.(1) . More spectra details are shown on S.M.sup (2020).Sec(VIII).
We now present the spectra of excitations of the quasi-hole ($N=3$) and quasi-particle ($N=5$) sectors as shown in Fig.(2) ($a$) and ($b$), respectively. Since each SO(5) IRREP contains SO(4) IRREPS $(j_{1},j_{2})_{so4}\oplus(j_{2},j_{1})_{so4}$ symmetrically, we only present the state in the sector $(\frac{1}{2},0)$ in terms of $(M_{J},M_{K})$ as a representative. As for the quasi-hole states of $N=3$, its ground states remains at zero energy with 20-fold degeneracy, lying in the SO(5) IRREP of $(3,0)_{so5}$, which is decomposed to SO(4) IRREPs of $(\frac{3}{2},0)_{so4}\oplus(0,\frac{3}{2})_{so4}\oplus(1,\frac{1}{2})_{so4}\oplus(\frac{1}{2},1)_{so4}$. In contrast, as for the quasi-particle states with $N=5$, there exist a finite excitation gap. The first excited states form the SO(5) IRREP of $(3,0)_{so5}$, which is the same as the quasi-hole case.
Early studies of FQHE debated whether the ground state should be an incompressible quantum liquid or a Wigner crystal Fukuyama et al. (1979); Yoshioka et al. (1983); Yoshioka and Lee (1983); Levesque et al. (1984); Yoshioka (1984). The pair distribution function is frequently employed to reveal the incompressible nature of quantum liquids Jain (2007). We calculate the probability $h(\theta)$ of two particles separated at a distance, which is a function of the angle $\theta$ between two unit vectors from the center to particles due to the rotation symmetry. It is defined as
$h(\theta)=\frac{\int d^{3}n_{1}d^{3}n_{2}\langle\Psi_{0}\mid\rho(x_{a},n_{1})\rho(x_{a}^{\prime},n_{2})\mid\Psi_{0}\rangle}{\rho(x_{a})\rho(x^{\prime}_{a})}-1$ (16)
where $|\Psi_{0}\rangle$ is the ground state of wavefunction Eq.(13); $\cos\theta=x_{a}x_{a}^{\prime}$; $\rho(x_{a},n_{1})$ is density operator; $\rho(x_{a})=\int d^{3}n_{1}\langle\Psi_{0}|\rho(x_{a},n_{1})|\Psi_{0}\rangle$. The calculated behavior of $h(\theta)$ is shown in Fig.(3), which is analogous to that of Landau level on the $S^{2}$ manifold Haldane and Rezayi (1985); Girvin et al. (1986). As $\theta$ goes to 0, $h(\theta)$ approaches 1-1 due to the Fermi statistics. Its amplitude features a smooth decay to zero in the entire range of angle without pronounced peaks. This result implies the absence of crystalline order, and is consistent with the quantum liquid like ground state in 4D.
Figure 3: The pair distribution function of the 4D-determinant Laughlin state lacks a peak at long distance, indicating liquid-like behavior.
Now we discuss the construction of Jastrow-type Laughlin wavefunction on $S^{4}$ based on the $SO(5)$ singlet $\psi_{\alpha}(x_{1},n_{1})\mathcal{R}_{\alpha\beta}\psi_{\beta}(x_{2},n_{2})$, which is a generalization of the SU(2) singlet function $(u_{i}v_{j}-u_{j}v_{i})$ Haldane (1983). The explicit wavefunction is expressed with the assistance of $\mathcal{R}$,
$\Psi_{\text{Jastrow}}^{m}=\prod\limits_{1\leq k<l\leq N}\left(\psi_{\alpha}(x_{k,a},n_{k,i})\mathcal{R}_{\alpha\beta}\psi_{\beta}(x_{l,a},n_{l,i})\right)^{m},$ (17)
The power of the spinor $\psi$ of each particle appearing in Eq.(17) is equal to l=m(N1)l=m(N-1), hence, the filling should be $\nu=N/d(l,0)$ with $d(l,0)$ defined in Eq.(3) approaching $1/(m^{3}N^{2})\to 0$ as $N\to\infty$. The parent Hamiltonian to this wavefunction is constrained in channel $\bigoplus\limits_{n_{1}+n_{2}\geq l-m+1}(n_{1}+n_{2},n_{1}-n_{2})_{so5}$. There is no well-defined thermodynamic limit for the Jastrow-type wavefunction.
In conclusion, we have studied the 4D incompressible quantum liquid based on the Landau level under an SU(2) monopole gauge potential, specifically using the determinant-Laughlin and Jastrow wavefunctions. The parent Hamiltonian for these wavefunctions was constructed within the pseudo-potential formalism. The numerical calculations show that the unique ground state is an SO(5) singlet. In addition, quasi-hole states remain at zero energy. In contrast, quasi-particle excitations are separated by a finite gap. The liquid-like behavior is further supported by the pair distribution function. However, direct numerical analysis of both determinant-type Laughlin and Jastrow wavefunctions for larger particle numbers is prohibitive due to the enormous size of the Hilbert space. For $N=10$, the next accessible system size, the Hilbert space has a dimension of about $10^{11}$.Nonetheless, the recent experimental demonstration of the few-electron Coulomb liquid Mukherjee et al. (2022); Lunt et al. (2024); Shaju et al. (2025) provides a promising platform for the experimental validation of our theoretical framework. Another intriguing avenue is to explore the statistical properties of excitations in the 4D FQHE, particularly investigating loop or membrane statistics in higher dimensions Congjun Wu and Zhang (2010); Wang and Levin (2014); Chen et al. (2023b); Bernevig et al. (2002); Wang et al. (2019); Huang et al. (2024); Huxford et al. (2024); Agarwal et al. (2025). Furthermore, studying the entanglement spectrum Kitaev and Preskill (2006); Levin and Wen (2006); Li and Haldane (2008) would offer valuable insights into these states and warrants further investigation.
Acknowledgment.— We are grateful to Kun Yang, Hua Chen, Chen Lu, L.Q. Chen, Zhiming Pan, JianKeng Yuan, Jianyu Wang and Ziyuan Zeng for stimulating discussions. C.W. is supported by the National Natural Science Foundation of China under the Grants No. 12234016 and No. 12174317. W.Z. was supported by National Science Foundation of China under Grants No. 12474144. This work has been supported by the New Cornerstone Science Foundation.